THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 117 — 1 June 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 117^{th} issue of the Magellanic Clouds Newsletter. It features exciting new results on massive stars, pulsars, the structure of the Magellanic Clouds and dust production within them, as well as other no-less interesting work. If you want your work advertised on the cover just e-mail a nice picture to mcnews@astro.keele.ac.uk. The next issue is planned to be distributed on the 1^{st} of August 2012.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

The VLT–FLAMES Tarantula Survey V. The peculiar B[e]-like supergiant, VFTS 698, in 30 Doradus

P.R. Dunstall^{1}, M. Fraser^{1}, J.S. Clark^{2}, P.A. Crowther^{3}, P.L. Dufton^{1}, C.J. Evans^{4}, D.J. Lennon^{5}, I. Soszyński^{6}, W.D. Taylor^{7} and J.S. Vink^{8}

1Department of Physics & Astronomy, The Queen’s University of Belfast, BT7 1NN, Northern Ireland, UK 2Department of Physics and Astronomy, The Open University, Walton Hall, Milton Keynes, MK7 6AA, UK 3Department of Physics & Astronomy, University of Sheffield, Sheffield S3 7RH, UK 4UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 5ESA, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 6Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 7Institute for Astronomy, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 8Armagh Observatory, College Hill, Armagh BT61 9DG, Northern Ireland, UK

We present an analysis of a peculiar supergiant B-type star (VFTS 698/Melnick 2/Parker 1797) in the 30 Doradus region of the Large Magellanic Cloud which exhibits characteristics similar to the broad class of B[e] stars. We analyse optical spectra from the VLT–FLAMES survey, together with archival optical and infrared photometry and X-ray imaging to characterise the system. We find radial velocity variations of around 400 km s^{−1} in the high excitation Si iv, N iii and He ii spectra, and photometric variability of ∼ 0.6 mag with a period of 12.7 d. In addition, we detect long-term photometric variations of ∼ 0.25 mag, which may be due to a longer-term variability with a period of ∼ 400 d. We conclude that VFTS 698 is likely an interacting binary comprising an early B-type star secondary orbiting a veiled, more massive companion. Spectral evidence suggests a mid-to-late B-type primary, but this may originate from an optically-thick accretion disc directly surrounding the primary.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1201.6389

X-ray phase-resolved spectroscopy of PSRs B0531+21, B1509−58, and B0540−69 with RXTE

M.Y. Ge^{1}, F.J. Lu^{1}, J.L. Qu^{1}, S.J. Zheng^{1}, Y. Chen^{1} and D.W. Han^{1}

1Key Laboratory for Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, China

The Rossi X-Ray Timing Explorer (RXTE) has made hundreds of observations on three famous young pulsars (PSRs) B0531+21 (Crab), B1509−58, and B0540−69. Using the archive RXTE data, we have studied the phase-resolved spectral properties of these pulsars in details. The variation of the X-ray spectrum with phase of PSR B0531+21 is confirmed here much more precisely and more details are revealed than in the previous studies: The spectrum softens from the beginning of the first pulse, turns to harden right at the pulse peak and becomes the hardest at the bottom of the bridge, softens gradually until the second peak, and then softens rapidly. Different from the previous studies, we found that the spectrum of PSR B1509−58 is significantly harder in the center of the pulse, which is also in contrast to that of PSR B0531+21. The variation of the X-ray spectrum of PSR B0540−69 seems similar to that of PSR B1509−58, but with a lower significance. Using about 10 years of data span, we also studied the real time evolution of the spectra of these pulsars, and no significant evolution has been detected. We discuss the constraints of these results on theoretical models of pulsar X-ray emission.

Published in The Astrophysical Journal Supplement Series Available from arXiv:1204.2199

First evidence of globular cluster formation from the ejecta of prompt type Ia supernovae

Takuji Tsujimoto^{1} and Kenji Bekki^{2}

1National Astronomical Observatory of Japan, Japan 2ICRAR, Australia

Recent spectroscopic observations of globular clusters (GCs) in the Large Magellanic Cloud (LMC) have discovered that one of the intermediate-age GC, NGC 1718 with [Fe/H] = −0.7 has an extremely low [Mg/Fe] ratio of ∼ −0.9. We propose that NGC 1718 was formed from the ejecta of type Ia supernovae (SNe Ia) mixed with very metal-poor ([Fe/H] < −1.3) gas about 2 Gyr ago. The proposed scenario is shown to be consistent with the observed abundances of Fe-group elements such as Cr, Mn, and Ni. In addition, compelling evidence for asymptotic giant branch stars playing a role in chemical enrichment during this GC formation is found. We suggest that the origin of the metal-poor gas is closely associated with the efficient gas-transfer from the outer gas disk of the Small Magellanic Cloud to the LMC disk. We anticipate that the outer part of the LMC disk contains field stars exhibiting significantly low [Mg/Fe] ratios, formed through the same process as NGC 1718.

Accepted for publication in ApJ Letters Available from arXiv:1204.4213

The near-infrared detection of PSR B0540−69 and its nebula

R.P. Mignani^{1,2}, A. De Luca^{3,4,5}, W. Hummel^{6}, A. Zajczyk^{7,8}, B. Rudak^{7,9}, G.Kanbach^{10} and A. Słowikowska^{2}

1Mullard Space Science Laboratory, University College London, UK 2Kepler Institute of Astronomy, University of Zielona Góra, Poland 3INAF — Istituto di Astrofisica Spaziale e Fisica Cosmica, Milano, Italy 4IUSS — Istituto Universitario di Studi Superiori, Pavia, Italy 5INFN — Istituto Nazionale di Fisica Nucleare, Pavia, Italy 6European Southern Observatory, Garching, Germany 7Nicolaus Copernicus Astronomical Center, Torun, Poland 8LUPM, Université Montpellier 2, France 9KAA UMK, Torun, Poland 10Max-Planck Institut für Extraterrestrische Physik, Garching, Germany

The ∼ 1700 year old PSR B0540−69 in the Large Magellanic Clouds (LMC) is considered the twin of the Crab pulsar because of its similar spin parameters, magnetic field, and energetics. PSR B0540−69 (V ∼ 22.5 mag) is also one of the very few pulsars for which both optical pulsations and polarised emission have been measured. Its optical spectrum is fit by a power-law, ascribed to synchrotron radiation, like for the young Crab and Vela pulsars. At variance with them, however, a double break is required to join the X-ray and optical power-law spectra, with the first one possibly occurring in the near ultraviolet (nUV). Near-infrared (nIR) observations, never performed for PSR B0540−69, are crucial to determine whether the optical power-law spectrum extends to longer wavelengths or a new break occurs, like it happens for both the Crab and Vela pulsars in the mIR, hinting at an even more complex particle energy and density distribution in the pulsar magnetosphere. We observed PSR B0540−69 in the J, H, and K_{s} bands with the Very Large Telescope (VLT) to detect it, for the first time, in the nIR and characterise its optical-to-nIR spectrum. To disentangle the pulsar emission from that of its pulsar wind nebula (PWN), we obtained high-spatial resolution adaptive optics images with the NAOS–CONICA instrument (NACO). We could clearly identify PSR B0540−69 in our J, H, and K_{s}-band images and measure its flux (J = 20.14, H = 19.33, K_{s} = 18.55 mag, with an overall error of ±0.1 magnitudes in each band). The joint fit to the available optical and nIR photometry with a power-law spectrum F_{ν} ∝ ν^{−α} gives a spectral index α = 0.70 ± 0.04, slightly more precise than measured in the optical only. This suggests that there is no spectral break between the optical and the nIR. We also detected, for the first time, the PSR B0540−69 PWN in the nIR. The comparison between our NACO images and Hubble Space Telescope (HST) optical ones does not reveal any apparent difference in the PWN morphology as a function of wavelength. The PWN optical-to-nIR spectrum is also fit by a single power-law, with spectral index α = 0.56 ± 0.03, slightly flatter than the pulsar’s. Using NACO at the VLT, we obtained the first detection of PSR B0540−69 and its PWN in the nIR. Due to the small angular scale of the PWN (∼ 4′′) only the spatial resolution of the James Webb Space Telescope (JWST) will make it possible to extend the study of the pulsar and PWN spectrum towards the mid-IR.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1204.6655

The VPOS: a vast polar structure of satellite galaxies, globular clusters and streams around the Milky Way

Marcel S. Pawlowski^{1}, Jan Pflamm-Altenburg^{1} and Pavel Kroupa^{1}

1Argelander-Institute for Astronomy, University of Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany

It has been known for a long time that the satellite galaxies of the Milky Way (MW) show a significant amount of phase-space correlation, they are distributed in a highly inclined Disc of Satellites (DoS). We have extended the previous studies on the DoS by analysing for the first time the orientations of streams of stars and gas, and the distributions of globular clusters within the halo of the MW. It is shown that the spatial distribution of MW globular clusters classified as young halo clusters (YH GC) is very similar to the DoS, while 7 of the 14 analysed streams align with the DoS. The probability to find the observed clustering of streams is only 0.3 per cent when assuming isotropy. The MW thus is surrounded by a vast polar structure (VPOS) of subsystems (satellite galaxies, globular clusters and streams), spreading from Galactocentric distances as small as 10 kpc out to 250 kpc. These findings demonstrate that a near-isotropic infall of cosmological sub-structure components onto the MW is essentially ruled out because a large number of infalling objects would have had to be highly correlated, to a degree not natural for dark matter sub-structures. The majority of satellites, streams and YH GCs had to be formed as a correlated population. This is possible in tidal tails consisting of material expelled from interacting galaxies. We discuss the tidal scenario for the formation of the VPOS, including successes and possible challenges. The potential consequences of the MW satellites being tidal dwarf galaxies are severe. If all the satellite galaxies and YH GCs have been formed in an encounter between the young MW and another gas-rich galaxy about 10–11 Gyr ago, then the MW does not have any luminous dark-matter substructures and the missing satellites problem becomes a catastrophic failure of the standard cosmological model.

Accepted for publication in MNRAS Available from arXiv:1204.5176

The XMM–Newton survey of the Small Magellanic Cloud: XMMU J010633.1−731543 and XMMU J010743.1−715953, two new

Be/X-ray binary systems

M.J. Coe^{1}, F. Haberl^{2}, R. Sturm^{2}, E.S. Bartlett^{1}, D. Hatzidimitriou^{3}, L.J. Townsend^{1}, A. Udalski^{4}, S. Mereghetti^{5} and M. Filipović^{6}

1University of Southampton, UK 2MPE, Germany 3University of Athens, Greece 4Warsaw University Observatory, Poland 5INAF, Milano, Italy 6University of Western Sydney, Australia

In the course of the XMM–Newton survey of the Small Magellanic Cloud (SMC), two new bright X-ray sources were discovered exhibiting the spectral characteris- tics of High Mass X-ray Binaries – but revealing only weak evidence for pulsations in just one of the objects (at 153s in XMMU J010743.1−715953). The accurate X- ray source locations permit the identification of these X-ray source with Be stars, thereby strongly suggesting these systems are new Be/X-ray binaries. From blue spectra the proposed classification for XMMU J010633.1−731543 is B0.5–1Ve and for XMMU J010743.1−715953 it is B2IV–Ve.

Accepted for publication in MNRAS Available from arXiv:1204.5903

Stellar mass and age determinations. I. Grids of stellar models from Z = 0.006 to 0.04 and M = 0.5 to 3.5 M_{⊙}

N. Mowlavi^{1,2}, P. Eggenberger^{1}, G. Meynet^{1}, S. Ekström^{1}, C. Georgy^{3}, A. Maeder^{1}, C. Charbonnel^{1} and L. Eyer^{1}

1Geneva Observatory, Switzerland 2ISDC, Geneva Observatory, Switzerland 3Centre de Recherche Astrophysique, École Normale Supérieure de Lyon, France

We present dense grids of stellar models suitable for comparison with observable quantities measured with great precision, such as those derived from binary systems or planet-hosting stars. We computed new Geneva models without rotation at metallicities Z = 0.006, 0.01, 0.014, 0.02, 0.03 and 0.04 (i.e. [Fe/H] from −0.33 to +0.54) and with mass in small steps from 0.5 to 3.5 M_{⊙}. Great care was taken in the procedure for interpolating between tracks in order to compute isochrones. Several properties of our grids are presented as a function of stellar mass and metallicity. Those include surface properties in the Hertzsprung–Russell diagram, internal properties including mean stellar density, sizes of the convective cores, and global asteroseismic properties. We checked our interpolation procedure and compared interpolated tracks with computed tracks. The deviations are less than 1% in radius and effective temperatures for most of the cases considered. We also checked that the present isochrones provide nice fits to four couples of observed detached binaries and to the observed sequences of the open clusters NGC 3532 and M 67. Including atomic diffusion in our models with M < 1.1 M_{⊙} leads to variations in the surface abundances that should be taken into account when comparing with observational data of stars with measured metallicities. For that purpose, iso-Z_{surf} lines are computed. These can be requested for download from a dedicated web page together with tracks at masses and metallicities within the limits covered by the grids. The validity of the relations linking Z and [Fe/H] is also re-assessed in light of the surface abundance variations in low-mass stars.

Published in A&A Available from arXiv:1201.3628 and from http://www.aanda.org/index.php?option=com article&access=doi&doi=10.1051/0004-6361/201117749&Itemid=129

The mass-loss return from evolved stars to the Large Magellanic Cloud VI: Luminosities and mass-loss rates on population scales

D. Riebel^{1}, S. Srinivasan^{2}, B. Sargent^{3,4} and M. Meixner^{4,1}

1Department of Physics and Astronomy, The Johns Hopkins University, 3400 North Charles St. Baltimore, MD 21218, USA 2UPMC–CNRS UMR7095, Institut d’Astrophysique de Paris, F-75014 Paris, France 3Center for Imaging Science and Laboratory for Multiwavelength Astrophysics, Rochester Institute of Technology, 54 Lomb Memorial Drive, Rochester, NY 14623, USA 4Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We present results from the first application of the Grid of Red Supergiant and Asymptotic Giant Branch ModelS (GRAMS) model grid to the entire evolved stellar population of the Large Magellanic Cloud (LMC). GRAMS is a pre-computed grid of 80,843 radiative transfer (RT) models of evolved stars and circumstellar dust shells composed of either silicate or carbonaceous dust. We fit GRAMS models to ∼ 30, 000 Asymptotic Giant Branch (AGB) and Red Supergiant (RSG) stars in the LMC, using 12 bands of photometry from the optical to the mid-infrared. Our published dataset consists of thousands of evolved stars with individually determined evolutionary parameters such as luminosity and mass-loss rate. The GRAMS grid has a greater than 80% accuracy rate discriminating between oxygenand carbon-rich chemistry. The global dust injection rate to the interstellar medium (ISM) of the LMC from RSGs and AGB stars is on the order of 1.5 × 10^{−5} M_{⊙} yr^{−1}, equivalent to a total mass injection rate (including the gas) into the ISM of ∼ 5 × 10^{−3} M_{⊙} yr^{−1}. Carbon stars inject two and a half times as much dust into the ISM as do O-rich AGB stars, but the same amount of mass. We determine a bolometric correction factor for C-rich AGB stars in the K band as a function of J–K color, BC(K) = −0.40(J −K)^{2} +1.83(J −K)+1.29. We determine several IR color proxies for the dust mass-loss rate (MLR) from C-rich AGB stars, such as log(MLR) = −18.90/((K − [8.0]) + 3.37) − 5.93. We find that a larger fraction of AGB stars exhibiting the ‘long-secondary period’ phenomenon are O-rich than stars dominated by radial pulsations, and AGB stars without detectable mass loss do not appear on either the first-overtone or fundamental-mode pulsation sequences.

Accepted for publication in The Astrophysical Journal Available from arXiv:1205.0280

A new observational tracer for high-density disc-like structures around B[e] supergiants

Anna Aret^{1,2}, Michaela Kraus^{1}, Maria Florencia Muratore^{3,4} and Marcelo Borges Fernandes^{5}

1Astronomický ústav, Akademie věd České republiky, Fričova 298, 251 65 Ondřejov, Czech Republic 2Tartu Observatory, 61602, Tõravere, Tartumaa, Estonia 3Departamento de Espectroscopía Estelar, Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata,Paseo del Bosque s/n, B1900FWA, La Plata, Argentina 4Instituto de Astrofísica de La Plata, CCT La Plata, CONICET–UNLP, Paseo del Bosque s/n, B1900FWA, La Plata, Argentina 5Observatório Nacional, Rua General José Cristino 77, 20921-400 São Cristovão, Rio de Janeiro, Brazil

The disc formation mechanism of B[e] supergiants is one of the puzzling phenomena in massive star evolution. Rapid stellar rotation seems to play an important role for the non-spherically symmetric mass-loss leading to a high-density disc or ring-like structure of neutral material around these massive and luminous objects. The radial density and temperature structure as well as the kinematics within this high-density material are, however, not well studied. Based on high-resolution optical spectra of a sample of B[e] supergiants in the Magellanic Clouds we especially searched for tracers of the kinematics within their discs. Besides the well-known [O i] lines, we discovered the [Ca ii] λλ7291, 7324 lines that can be used as a complementary set of disc tracers. We find that these lines originate from very high-density regions, located closer to the star than the [O i] λ5577 line-forming region. The line profiles of both the [O i] and the [Ca ii] lines indicate that the discs or rings of high-density material are in Keplerian rotation. We estimate plausible ranges of disc inclination angles for the sample of B[e] supergiants and suggest that the star LHA 120-S 22 might have a spiral arm rather than a disc.

Accepted for publication in MNRAS Available from arXiv:1203.6808

A rare early-type star revealed in the Wing of the Small Magellanic Cloud

C.J. Evans ^{1}, R. Hainich^{2}, L.M. Oskinova^{2}, J.S Gallagher^{3}, Y.-H. Chu^{4}, R.A. Gruendl^{4}, W.-R. Hamann^{2}, V. Hénault-Brunet^{5} and H. Todt^{2}

1UK Astronomy Technology Centre, Edinburgh, Scotland 2University of Potsdam, Germany 3University of Wisconsin–Madison, USA 4University of Illinois, USA 5University of Edinburgh, Scotland

Sk 183 is the visually-brightest star in the N 90 nebula, a young star-forming region in the Wing of the Small Magellanic Cloud (SMC). We present new optical spectroscopy from the Very Large Telescope which reveals Sk 183 to be one of the most massive O-type stars in the SMC. Classified as an O3-type dwarf on the basis of its nitrogen spectrum, the star also displays broadened He i absorption which suggests a later type. We propose that Sk 183 has a composite spectrum and that it is similar to another star in the SMC, MPG 324. This brings the number of rare O2- and O3-type stars known in the whole of the SMC to a mere three. We estimate physical parameters for Sk 183 from analysis of its spectrum. For a single-star model, we estimate an effective temperature of 46 ± 2 kK, a low mass-loss rate of ∼ 10^{−7} M_{⊙} yr^{−1}, and a spectroscopic mass of 46_{−8}^{+9} M_{⊙} (for an adopted distance modulus of 18.7 mag to the young population in the SMC Wing). An illustrative binary model requires a slightly hotter temperature (∼ 47.5 kK) for the primary component. In either scenario, Sk 183 is the earliest-type star known in N 90 and will therefore be the dominant source of hydrogen-ionising photons. This suggests Sk 183 is the primary influence on the star formation along the inner edge of the nebula.

Accepted for publication in ApJ Available from arXiv:1205.2387

The nature of dust in compact Galactic planetary nebulae from Spitzer spectra

L. Stanghellini^{1}, D.A. García-Hernández^{2}, P. García-Lario^{3}, J.E. Davies^{4}, R.A. Shaw^{1}, E. Villaver^{5}, A. Manchado^{2} and J.V. Perea-Calderón^{6}

1NOAO, USA 2IAC–ULL, Spain 3ESAC–ESA, Spain 4Spitzer Science Center–CIT, USA 5UAM, Spain 6INSA–ESAC, Spain

We present the Spitzer/IRS spectra of 157 compact Galactic planetary nebulae (PNe). These young PNe provide insight on the effects of dust in early post-AGB evolution, before much of the dust is altered or destroyed by the hardening stellar radiation field. Most of the selected targets have PN-type IRS spectra, while a few turned out to be misclassified stars. We inspected the group properties of the PN spectra and classified them based on the different dust classes (featureless, or F; carbon-rich dust, or CRD; oxygen-rich dust, or ORD; mixed-chemistry dust, or MCD) and subclasses (aromatic and aliphatic; crystalline and amorphous). All PNe are characterized by dust continuum and more than 80% of the sample shows solid state features above the continuum, in contrast with the Magellanic Cloud sample where only ∼ 40% of the entire sample displays solid state features; this is an indication of the strong link between dust properties and metallicity. The Galactic PNe that show solid state features are almost equally divided among the CRD, ORD, and MCD. We analyzed dust properties together with other PN properties and found that (i) there is an enhancement of MCD PNe toward the Galactic center, in agreement with studies of Galactic bulge PNe; (ii) CRD PNe could be seen as defining an evolutionary sequence, contrary to the ORD and MCD PNe, which are scattered in all evolutionary diagrams; (iii) carbon-rich and oxygen-rich grains retain different equilibrium temperatures, as expected from models; (iv) ORD PNe are highly asymmetric, i.e. bipolar or bipolar-core, and CRD PNe highly symmetric, i.e. round or elliptical; point-symmetry is statistically more common in MCD than in other dust class PNe. By comparing the sample of this paper to that of Magellanic Cloud PNe we find that the latter sample does not include MCD PNe, and the other dust classes are differently populated, with continuity of the fraction of F, CRD, ORD, and MCD population from high to low metallicity environments. We also find similar sequences for CRD PNe in the Galactic disk and the Magellanic Clouds, except that the Magellanic Cloud PNe seem to attain higher dust temperatures at similar evolutionary stages, in agreement with the observational findings of smaller dust grains (i.e. lower radiation efficiency) in low metallicity interstellar environments.

Accepted for publication in ApJ Available from arXiv:1205.3829 and from http://rialto.ll.iac.es/folleto/research/preprints/?c=view&pre id=12018

The period–luminosity relation of red supergiant stars in the Small Magellanic Cloud

Ming Yang^{1} and Biwei Jiang^{1}

1Department of Astronomy, Beijing Normal University, China

The characteristics of light variation of RSGs in SMC are analyzed based on the nearly 8–10 year long data collected by the ASAS and MACHO projects. The identified 126 RSGs are classified into five categories accordingly: 20 with poor photometry, 55 with no reliable period, 6 with semi-regular variation, 15 with Long Secondary Period (LSP) and distinguishable short period and 30 with only LSP. For the semi-regular variables and the LSP variables with distinguishable short period, the K_{s}-band period–luminosity (P–L) relation is analyzed and compared with that of the Galaxy, LMC and M 33. It is found that the RSGs in these galaxies obey similar P–L relation except the Galaxy. In addition, the P–L relations in the infrared bands, namely the 2MASS JHK_{s}, Spitzer/IRAC and Spitzer/MIPS 24 μm bands, are derived with high reliability. The best P–L relation occurs in the Spitzer/IRAC [3.6] and [4.5] bands. Based on the comparison with the theoretical calculation of the P–L relation, the mode of pulsation of RSGs in SMC is suggested to be the first overtone radial mode.

Accepted for publication in The Astrophysical Journal Available from arXiv:1205.1275

Constraints on the Magellanic Clouds’ interaction from the distribution of OB stars and the kinematics of giants

Dana I. Casetti-Dinescu^{1}, Katherine Vieira^{2}, Terrence M. Girard^{1} and William F. van Altena^{1}

1Yale University, USA 2Centro de Investigaciones de Astronomía, Venezuela

Young, OB-type candidates are identified in a ∼ 7900 sq. deg. region encompassing the Large and Small Magellanic Clouds (LMC/SMC) periphery, the Bridge, part of the Magellanic Stream (MS) and Leading Arm (LA). Selection is based on UV, optical and IR photometry from existing large-area surveys and proper motions from the Southern Proper Motion 4 catalog (SPM4). The spatial distribution of these young star candidates shows 1) a well-populated SMC wing which continues westward with two branches partially surrounding the SMC, 2) a rather narrow path from the SMC wing eastward toward the LMC which is offset by 1 to 2 degrees from the high-density H i ridge in the Bridge,

3) a well-populated periphery of the LMC dominated by clumps of stars at the ends of the LMC bar and 4) a few scattered candidates in the MS and two overdensities in the LA regions above and below the Galactic plane. Additionally, a proper-motion analysis is made of a radial-velocity selected sample of red giants and supergiants in the LMC, previously shown to be a kinematically and chemically distinct subgroup, most likely captured from the SMC. SPM4 proper motions of these stars also indicate they are distinct from the LMC population. The observational results presented here, combined with the known orbits of the Clouds, and other aspects of the LMC morphology, suggest an off-center, moderate to highly-inclined collision between the SMC and the LMC’s disk that took place between 100 and 200 Myr ago.

Accepted for publication in Astrophysical Journal Available from arXiv:1205.0989 and from http://www.astro.yale.edu/dana/Clouds.pdf

The origin of the microlensing events observed towards the LMC and the stellar counterpart of the Magellanic Stream

Gurtina Besla^{1}, Lars Hernquist^{2} and Abraham Loeb^{2}

1Columbia University, USA 2Harvard-Smithsonian Center for Astrophysics, USA

We introduce a novel theory to explain the long-standing puzzle of the nature of the microlensing events reported towards the Large Magellanic Cloud (LMC) by the MACHO and OGLE collaborations. We propose that a population of tidally stripped stars from the Small Magellanic Clouds (SMC) located ∼ 4–10 kpc behind a lensing population of LMC disk stars can naturally explain the observed event durations, event frequency and spatial distribution of the reported events. These results favor a scenario for the interaction history of the Magellanic Clouds, wherein the Clouds are on their first infall towards the Milky Way and the SMC has recently collided with the LMC 100–300 Myr ago, leading to a large number of faint sources distributed non-uniformly behind the LMC disk. Owing to the tidal nature of the source population, the sources exhibit a range of distances and velocities with respect to the LMC lenses, naturally explaining the observed range of event durations (30–220 days). Differences in the event frequencies reported by the OGLE and MACHO surveys appear to be naturally accounted for by their different detection efficiencies and sensitivity to faint sources. Assuming a detection efficiency of 30–50% we find event frequencies of Γ ∼ 1–2 yr^{−1} in the central regions of the LMC disk; comparable to the observed rate for the MACHO survey, Γ ∼ 2 yr^{−1}. A lower detection efficiency of 10% yields an event frequency of Γ ∼ 0.46 yr^{−1} across a larger area of the LMC disk; comparable to that reported by the less sensitive OGLE survey, Γ ∼ 0.33 yr^{−1}. In contrast to self-lensing models, microlensing events are also expected to occur in fields off the LMC’s stellar bar since the stellar debris is not expected to be concentrated in the bar region. This scenario leads to a number of observational tests: the sources are low-metallicity SMC stars, they exhibit high velocities relative to LMC disk stars that may be detectable via proper motion studies, and, most notably, there should exist a stellar counterpart to the gaseous Magellanic Stream and Magellanic Bridge with a V-band surface brightness of > 34 mag arcsec^{−2}.

Submitted to MNRAS Available from arXiv:1205.4724

Integrated J- and H-band spectra of globular clusters in the LMC: implications for stellar population models and galaxy age dating

Mariya Lyubenova^{1}, Harald Kuntschner^{2}, Marina Rejkuba^{2}, David R. Silva^{3}, Markus Kissler-Patig^{2} and Lowell E. Tacconi-Garman^{2}

1Max Planck Institute for Astronomy, Königstuhl 17, D-69117 Heidelberg, Germany 2European Southern Observatory, Karl-Schwarzschild-Str. 2, D-85748 Garching bei München, Germany 3National Optical Astronomy Observatory, 950 North Cherry Ave., Tucson, AZ, 85719 USA

(Abridged) The rest-frame near-IR spectra of intermediate age (1–2 Gyr) stellar populations are dominated by carbon based absorption features offering a wealth of information. Yet, spectral libraries that include the near-IR wavelength range do not sample a sufficiently broad range of ages and metallicities to allow for accurate calibration of stellar population models and thus the interpretation of the observations. In this paper we investigate the integrated J- and H-band spectra of six intermediate age (1–3 Gyr) and old (> 10 Gyr) globular clusters in the Large Magellanic Cloud, using observations obtained with the SINFONI IFU at the VLT. H-band C_{2} and K-band ^{12}CO(2–0) feature strengths are compared to the models of Maraston (2005). C_{2} is reasonably well reproduced by the models at all ages, while 12CO(2–0) shows good agreement for older (age > 2 Gyr) populations, but the younger (1.3 Gyr) globular clusters do not follow the models. We argue that this is due to the fact that the empirical calibration of the models relies on only a few Milky Way carbon star spectra, which show different ^{12}CO(2–0) index strengths than the LMC stars. The C_{2} absorption feature strength correlates strongly with age. It is present essentially only in populations that have 1–2 Gyr old stars, while its value is consistent with zero for older populations. The distinct spectral energy distribution observed for the intermediate age globular clusters in the J- and H-bands agrees well with the model predictions of Maraston for the contribution from the thermally pulsing asymptotic giant branch phase (TP-AGB). We show that the H-band C_{2} absorption feature and the J-, H-band spectral shape can be used as an age indicator for intermediate age stellar populations in integrated spectra of star clusters and galaxies.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1205.5269